OpenStudy (anonymous):

find the slope (-3,6) (6,-6)

OpenStudy (anonymous):

Use the slope formula: m= y2-y1 / x2- x1 y2= -6 y1= 6 x2= 6 x1=-3

OpenStudy (anonymous):

m=-6-6/6--3 the double negative you would add -12/9=-1.33

OpenStudy (anonymous):

there you got it (: